40.0 s after fring. What are To start on a mountain slope, an artiiery shell is fired with an initial velocity of 350 m/s at 49,0 above the and y coordinates of the shell where it explodes, relative to its fring point 12378 35 × Use the constant equations separately for the x components and for the y cemponents of the motion. m Need Help?
Answer #1

Vx = horizontal velocity = 350 cos 49=229.62 m/s

X= 229.62 (40) = 9184.826 m

y = 350 sin ( 49) (40 ) - 1/2 (9.8) ( 40)^2

y = 10565.934 - 7840 =2725.934 m